# Closing the Gap on the Sample Complexity of 1-Identification

### v1

@@ C.2 Properties of Algorithm 2
C.3 Proof of Theorem 2 and 3
The full proof requires multiple lemmas to support, which motivates us to use figure 1 to summarize the overall ideas. An arrow from A to B means Lemma A is used in the proof of Lemma/Theorem B. All the Lemmas inside the orange box are the properties of Algorithm 2, which are mainly adapted from Li and Cheung (2025).
Figure 1: Overall Idea of Full Proof of Theorem 2 and 3
@@ C.3 Proof of Theorem 2 and 3
Before illustrating the proof of Theorem 2 and 3, we first introduce a Lemma for the random permutation, Line 1 of Algorithm 1.
Lemma 31.
@@
The second last line is by Lemma 18 and 19.
@@
The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 17, 31. For the term $\ddagger$, we have
@@
Consider the bracket with size $K$, denoting this bracket as $B$. Similar to the proof of positive case, we split $\tau(B)=\tau^{\text{ee}}(B)+\tau^{\text{et}}(B)$ and prove an upper bound for $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$ separately. From Lemma 14, 29 and Lemma 30, we know
@@ Appendix D Technical Lemma
$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq s\big)$
$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda s\big)$
$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda s\right)\right).$
@@
$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda s)}$
$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda s)}$
$=$ $\exp(-\frac{s^{2}}{2n\Delta^{2}}).$
@@
Lemma 35 (Adapted Lemma 3 in Jamieson et al. (2014), Lemma D.3 in Li and Cheung (2025)).
@@
Lemma 36 (Lemma D.2 in Li and Cheung (2025)).

### v2

@@ C.2 Properties of Algorithm 2
C.3 Sketch Proof of Theorem 2, 3
The full proof of Theorem 2, 3 requires multiple lemmas to support.
Sketch Proof of Theorem 2.
The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung [2025]. As in Li and Cheung [2025], denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase indices after which the exploration and exploitation concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
Given the first step, we are able to conclude $\Pr(\tau<+\infty)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$. The two cases are proved similarly. In Line 1 of Algorithm 1, we assign $\delta/(\lceil\log_{2}K\rceil+1)$ to each copy. We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. A union bound over all copies completes the proof. ∎
Sketch Proof of Theorem 3.
We consider negative and positive instances separately.
If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung [2025]. By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. For $j\in[m]$ and $a\in[K]$, define $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$. We show that for every $j\in[m]$, $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$. In addition, the constant hidden by the Big O notation is independent of $j$. For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, the smallest bracket index that contains at least one arm among the top-$j$ arms. As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$. Using $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ for each $b\in[\lceil\log_{2}K\rceil+1]$, it suffices to bound $O(\log K)\,\mathbb{E}_{\nu}[\tau(B_{b})\mathds{1}(b_{j}=b)]$.
Let $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ denote the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ respectively. For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period. It is evident that $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$. We assert that
Given (32)(33), we can sum up (32)(33) over $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$. Lemma 7(section C.4), 33, 34( section D) provides upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3. The remaining sketch proof reduces to (32)(33) for each $b\in[\lceil\log_{2}K\rceil+1]$.
To prove (32), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$. Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 18 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (32).
To prove (33), we first show for each $a\in B_{b}$, we have
The intuition is as follows. Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is of order $O(1)\tilde{T}_{j}(a;L_{b}^{\text{pos}})(C+1)^{2}$, copy $\text{alg}_{b}$ can assert(with some probability, see Lemma 18, 25) either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$. In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period. By the facts that elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (34) over $a\in B_{b}$ and derive
Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, (33) is shown. ∎
C.4 Proof of Theorem 2 and 3
@@
In this subsection, We present the rigorous proof of Theorem 2 and 3 and the required lemmas. Before illustrating the proof of Theorem 2 and 3, we first prove Lemma 7 for the random permutation, Line 1 of Algorithm 1.
Lemma (restatement of Lemma 7).
@@
The second last line is by Lemma 19 and 20.
@@
The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 18, 7. For the term $\ddagger$, we have
@@
Consider the bracket with size $K$, denoting this bracket as $B$. Similar to the proof of positive case, we split $\tau(B)=\tau^{\text{ee}}(B)+\tau^{\text{et}}(B)$ and prove an upper bound for $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$ separately. From Lemma 15, 30 and Lemma 31, we know
@@ Appendix D Technical Lemma
$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq z\big)$
$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda z\big)$
$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda z\right)\right).$
@@
$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda z)}$
$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda z)}$
$=$ $\exp(-\frac{z^{2}}{2n\Delta^{2}}).$
@@
Lemma 35 (Adapted Lemma 3 in Jamieson et al. [2014], Lemma D.3 in Li and Cheung [2025]).
@@
Lemma 36 (Lemma D.2 in Li and Cheung [2025]).